OpenStudy (anonymous):

in rwacc and beta levered computations should I add the MV of convertibles and MV of option claims on equity to market value of equity?

OpenStudy (anonymous):

\[\leftarrow \int\limits_{?}^{?} \cot \cot \cot \cot \exp \cos^{-1} \tan \sin \sin \kappa \Pi \alpha \kappa \kappa \pi \phi \pi \rho \Psi \Psi \lambda \Lambda \Psi \nu \omega \Lambda \pi \pi \iota\]no

OpenStudy (anonymous):

You should add them because market value of equity computed from stock prices is the common equity value. You need to have to have total equity , which includes preferred shares, options of convertibles, and employee options. Refer to Damodaran quizzes where he clearly in one of them he adds equity portion of convertibles to market value of equity